A driver has to travel 40 miles at an average speed of 40 miles an hour. She maintains this average for 20 miles but for 10 miles she drops down maintained average of 40 miles an hour. In order to serve the remaining 10 miles and average 40 miles an hour she must drive a)60mi/h b)80 mi/h c)90 mi/h
1
Expert's answer
2018-10-02T08:40:08-0400
20 miles she travel for 20mi/(40 mi/h)=0.5 h
10 next miles (10 mi)/(30 mi/h)=1/3 h
10 next miles (10 mi)/x=1-0.5-1/3
X=60 mi/h